ISSN 1201-561X

[figure: T; A; C]

THEORY AND APPLICATIONS OF CATEGORIES

Volume 32, 2017

A structure theorem for quasi-Hopf bimodule coalgebras Daniel Bulacu, 1-30

Simplicial Nerve of an A_{∞}-category Giovanni Faonte, 31-52

C-systems defined by universe categories: presheaves Vladimir Voevodsky, 53-112

The (Π, λ)-structures on the C-systems defined by universe categories Vladimir Voevodsky, 113-121

Bourn-normal monomorphisms in regular Mal’tsev categories Giuseppe Metere, 122-147

Homotopy theory for algebras over polynomial monads M. A. Batanin and C. Berger, 148-253

A note on injective hulls of posemigroups Changchun Xia, Shengwei Han, Bin Zhao, 254-257

Witt vectors and truncation posets Vigleik Angeltveit, 258-285

Classifying tangent structures using Weil algebras Poon Leung, 286-337

KZ-monadic categories and their logic Jiri Adamek and Lurdes Sousa, 338-379

Corelations are the prop for extraspecial commutative Frobenius monoids Brandon Coya and Brendan Fong, 380-395 Monoid-like definitions of cyclic operad Jovana Obradovic, 396-436

\begin{matrix}c \\ ◯ \end{matrix} The Editors of Theory and Applications of Categories 2017. Permission to copy for private use

granted.

Building a model category out of multiplier ideal sheaves Seunghun Lee, 437-487

A categorical approach to Picard-Vessiot theory Andreas Maurischat, 488-525

Epireflective subcategories and formal closure operators Mathieu Duckerts-Antoine, Marino Gran, and Zurab Janelidze, 526-546

Algebraic databases Patrick Schultz, David I. Spivak, Christina Vasilakopoulou, and Ryan Wisnesky, 547-619

The two out of three property in ind-categories and a convenient model category of spaces Ilan Barnea, 620-651

Frobenius algebras and homotopy fixed points of group actions on bicategories Jan Hesse, Christoph Schweigert, and Alessandro Valentino, 652-681

A presentation of bases for parametrized iterativity Jiri Adamek, Stefan Milius and Jiri Velebil, 682-718

Closure operators in abelian categories and spectral spaces Abhishek Banerjee, 719-735

Lax distributive laws for topology, II Hongliang Lai, Lili Shen and Walter Tholen, 736-768

A property of effectivization and its uses in categorical logic Vasileios Aravantinos-Sotiropoulos and Panagis Karazeris, 769-779

Levi categories and graphs of groups Mark V. Lawson and Alistair R. Wallis, 780-802

Metric spaces and SDG Anders Kock, 803-822

A unifying approach to the acyclic models method and other lifting lemmas Leonard Guetta, 823-834

Connections in tangent categories J.R.B. Cockett and G.S.H. Cruttwell, 835-888

Biextensions, bimonoidal functors, multilinear functor calculus, and categorical rings Ettore Aldrovandi, 889-969

Urysohn type lemma for groupoids Madalina Roxana Buneci, 970-994

A bicategory of decorated cospans Kenny Courser, 995-1027 The canonical 2-gerbe of a holomorphic vector bundle Markus Upmeier, 1028-1049

Classifiers for monad morphisms and adjunction morphisms Dimitri Zaganidis, 1050-1097

Free objects over posemigroups in the category PoSgr_{v} Shengwei Han, Changchun Xia, Xiaojuan Gu, 1098-1106

Combinatorics of past-similarity in higher dimensional transition systems Philippe Gaucher, 1107-1164

A Dold-Kan theorem for simplicial Lie algebras P. Carrasco and A.M. Cegarra, 1165-1212

Localization of enriched categories and cubical sets Tyler Lawson, 1213-1221

Coexponentiability and Projectivity: Rigs, Rings, and Quantales S.B. Niefield and R.J. Wood, 1222-1228

Hopf polyads, Hopf categories and Hopf group monoids viewed as Hopf monads Gabriella Böhm, 1229-1257

Fibered multiderivators and (co)homological descent Fritz Hörmann, 1258-1362

Kan’s combinatorial spectra and their sheaves revisited Ruian Chen, Igor Kriz and Ales Pultr, 1363-1396

Operads and phylogenetic trees John C. Baez and Nina Otter, 1396-1453

Topological properties of non-Archimedean approach spaces Eva Colebunders and Karen Van Opdenbosch, 1454-1484

Mal’tsev objects, R_{1}-spaces and ultrametric spaces Thomas Weighill, 1485-1500

Action representability via spans and prefibrations J. R. A. Gray, 1501-1521

Approximate categorical structures Abdelkrim Aliouche and Carlos Simpson, 1522-1562

Stability properties characterising n-permutable categories Pierre-Alain Jacqmin, Diana Rodelo, 1563-1587

A characterization of the Huq commutator Vaino Tuhafeni Shaumbwa, 1588-1600

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive by e-mail abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX 2ε is strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files from http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis: cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com Valeria de Paiva: Nuance Communications Inc: valeria.depaiva@gmail.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne: kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Edinburgh: Tom.Leinster@ed.ac.uk Ieke Moerdijk, Utrecht University: i.moerdijk@uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Ljubljana: Alex.Simpson@fmf.uni-lj.si James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: ross.street@mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca